SkyChefs, Inc., prepares in-flight meals for a number of major airlines. One of the company’s products is grilled salmon in dill sauce with baby new potatoes and spring vegetables. During the most recent week, the company prepared 7,200 of these meals using 2,800 direct labor-hours. The company paid its direct labor workers a total of $22,400 for this work, or $8.00 per hour.

According to the standard cost card for this meal, it should require 0.40 direct labor-hours at a cost of $7.50 per hour.

Required:

1. What is the standard labor-hours allowed (SH) to prepare 7,200 meals?

2. What is the standard labor cost allowed (SH × SR) to prepare 7,200 meals?

3. What is the labor spending variance?

4. What is the labor rate variance and the labor efficiency variance?

(For requirements 3 and 4, indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values. Do no round intermediate calculations.)
Answer #1

solution 1) Standard Labor hour allowed = 7200*0.40 = 2880 Hours 2) Standard Cost allowed = 2880*7.5 = $21600 3) Labor spendi